A 2.8 kg block of wood sits on a table. A 3.0 g bullet, fired horizontally at a speed of 550 m/s , goes completely through the block, emerging at a speed of 220 m/s .

What is the speed of the block immediately after the bullet exits? Express your answer to two significant figures and include the appropriate units.
Answer #1

here,

mass of wooden block , m1 = 2.8 kg

mass of bullet , m2 = 3 g = 0.003 kg

initial speed of bullet , u2 = 550 m/s

final speed of bullet , v2 = 220 m/s

let the final speed of block be v1

using conservation of momentum

m2 * u2 = m1 * v1 + m2 * v2

0.003 * 550 = 2.8 * v1 + 0.003 * 220

solving for v1

v1 = 0.35 m/s

the final speed of block after the collison is 0.35 m/s